The operations manager of a large production plant would like to estimate the


amount of time a worker takes to assemble a new electronic component. After


observing 120 of the workers assembling similar devices, the manager noticed


that the mean time was 16.2 minutes. Assuming that the population standard


deviation is 3.6 minutes;



c. Calculate the confidence level that would be required to have the mean


assembly time estimated up to ±15 seconds assuming the sample size


is 120 minutes.



1
Expert's answer
2021-12-13T16:07:52-0500

Given "n=120,\bar{x}=16.2\ min, \sigma=3.6\ min,"

"E=15\ sec=0.25\ min"


"CI=(\bar{x}-E, \bar{x}+E)"

"=(16.2-0.25,16.2+0.25)"

"=(15.95, 16.45)"


"E=z_c\times\dfrac{\sigma}{\sqrt{n}}"

"z_c=\dfrac{E(\sqrt{n})}{\sigma}"

"z_c=\dfrac{0.25(\sqrt{120})}{3.6}"

"z_c\approx0.7607"

"P(Z>0.7607)=1-P(Z\leq0.7607)"

"\approx1-0.77658\approx0.2234"

"2(0.2234)=0.4468"

"1-0.4468=0.5532"



"55.31"% confidence interval is "(15.95, 16.45)."


Therefore, based on the data provided, the 55.32 % confidence interval for the population mean is "15.95 < \mu < 16.45," which indicates that we are 

55.32% confident that the true population mean "\mu" is contained by the interval "(15.95, 16.45)."